Prepare a suppository containing 100 mg of phenobarbital (f = 0.81) using cocoa butter as the base.The weight of the pure cocoa butter suppository is 2.0 g. Because 100 mg of phenobarbi- tal is to be contained in an approximately 2.0-g suppository, it will be about 5% phenobarbital. What will be the total weight of each suppository?
Answer #1

answer) we have a formula for dosage replacement factor

f=(100(E-G)/(GX)+1

here E=weight of pure cocoa suppository

% of pehnobarbital =X

so we have all the values

0.81=100(2-G)/(G*5)+1

0.81=200-100G+5G/5G

4.05G=200-95G

99.05G=200

G=2.019 g

so the answer is 2.019 g ( if they want two significant figures then answer is 2.02 g but if its not mentioned how many significant figures then just write 2.019 g)